### v1

# Evolution Meets Diffusion: Efficient Neural Architecture Generation

Abstract
Neural Architecture Search (NAS) has gained widespread attention for its transformative potential in deep learning model design. However, the vast and complex search space of NAS leads to significant computational and time costs. Neural Architecture Generation (NAG) addresses this by reframing NAS as a generation problem, enabling the precise generation of optimal architectures for specific tasks. Despite its promise, mainstream methods like diffusion models face limitations in global search capabilities and are still hindered by high computational and time demands. To overcome these challenges, we propose Evolutionary Diffusion-based Neural Architecture Generation (EDNAG), a novel approach that achieves efficient and network-free architecture generation. EDNAG leverages evolutionary algorithms to simulate the denoising process in diffusion models, using fitness to guide the transition from random Gaussian distributions to optimal architecture distributions. This approach combines the optimization strengths of evolutionary strategies and the generation capabilities of diffusion models, enabling rapid and optimal architecture generation. Extensive experiments demonstrate that EDNAG achieves state-of-the-art (SOTA) performance in architecture optimization, with an improvement in accuracy of up to 10.45%. Furthermore, it eliminates the need for time-consuming training and boosts inference speed by an average of 50×, showcasing its exceptional efficiency.
1 Introduction
Figure 1: Overview of EDNAG. EDNAG generates neural architectures by simulating the denoising process in diffusion models. Unlike DiffusionNAG, which uses a trained noise prediction network to guide the denoising process, EDNAG employs fitness of architecture samples to derive the samples of next generation, thereby guiding the denoising process.
Neural Architecture Search (NAS) Dong et al. (2024); Kang et al. (2023) represents an effective approach for the automated design of deep neural network architectures. However, conventional NAS methods are limited by the significant computational costs associated with exploring vast and intricate search spaces. Recently, the paradigm of Neural Architecture Generation (NAG) Nasir et al. (2024); Lukasik et al. (2022) has revolutionized this optimization challenge by re-framing it from an optimization task into a generative task, thus mitigating the limitations imposed by large search spaces. The NAG framework leverages advanced generation techniques to directly produce optimized neural architectures that are finely tuned for specific downstream applications, offering a more scalable and resource-efficient solution.
Previous architecture generation techniques can be categorized into two primary classes: Variational Auto-Encoders (VAE)-based Lin et al. (2023) and Generative Adversarial Networks (GAN)-based Huang et al. (2023) methods. VAE-based methods train an encoder to map architecture features into latent space and a decoder to generate them from latent space, while GAN-based methods employ a generator to produce architectures with the reward from a discriminator. However, both VAE-based and GAN-based methods require the simultaneous training of two networks, leading to significant challenges in high computational complexity, training stability Brock (2018), and generation quality. Recently, diffusion-based methods have demonstrated superior performance by training a single network and generating architectures with multiple denoising steps, thereby improving stability and generation quality. For example, DiffusionNAG An et al. (2023) and DiNAS Asthana et al. (2024) employ a graph diffusion framework to generate architectures through iterative denoising steps, where each denoising iteration requires network inference to predict the noise to be eliminated. However, the generation process is inefficient due to the frequent network inferences and the multi-step denoising process, significantly slowing down the overall generation process. Moreover, when confronted with intricate and expansive search spaces, diffusion models often exhibit limitations in generating neural architectures with global optimality, as shown in Figure 2. Their naive diffusion models frequently converge to local optima, leading to the generation of suboptimal architectures. Therefore, diffusion-based methods still encounter substantial computational complexity during the training phase and exhibit suboptimality in the generation of neural architectures.
Figure 2: Each arrow in the figure represents a denoising step in diffusion models. DiffusionNAG simply adopts diffusion models, which tend to generate suboptimal architectures. In contrast, EDNAG leverages the global optimization capability of evolutionary algorithms to improve the generation process, making it more likely to produce globally optimal architectures.
To the end, we propose a novel and efficient Evolutionary Diffusion-based Neural Architecture Generation Framework, called EDNAG. The core of EDNAG lies in improving efficiency while improving generation quality. Specifically, we introduce evolutionary algorithms to simulate the denoising process in diffusion models. Furthermore, we propose a Fitness-guided Denoising (FD) strategy to produces a new generation of architectures from the previous ones in each denoising iteration. Each newly produced architecture can be conceptualized as a weighted summation of prior samples, with superior architectures assigned higher weights to exert a more prominent influence. Therefore, architecture samples will progressively evolve towards high-fitness subspaces through multiple denoising iterations with FD strategy, thereby generating high-performing and global-optimal neural architectures. In particular, EDNAG performs task-specific neural architecture generation with a neural predictor, serving as a dataset-aware fitness evaluator, to guide the conditional generation process with the FD strategy. Notably, the FD strategy eliminates the dependency on the score network in traditional diffusion models during the denoising process and achieves network-free neural architecture generation, therefore greatly improving generation efficiency.
We performed extensive experiments within NAS-Bench-101 Ying et al. (2019) on CIFAR10 and NAS-Bench-201 Dong and Yang (2020) across four task datasets. These experiments demonstrate that EDNAG not only achieves state-of-the-art (SOTA) performance in architecture generation with up to 10.45% improvement in accuracy, but also exhibits an average 50× speedup compared to previous baselines. Furthermore, we expanded the search space to TransNASBench-101 Duan et al. (2021) and MobileNetV3 Howard et al. (2019). The former search space involves seven distinct downstream computer vision tasks, while the latter features an extremely large search space. Experiments demonstrate that EDNAG consistently achieves state-of-the-art performance in architecture generation, highlighting its adaptability to larger search spaces and its transferability to diverse downstream tasks. Our code is available at https://github.com/cloud-zhoubingye/EvoDiff.git.
In summary, our contributions are as follows:
  • We propose a novel Evolutionary Diffusion-based framework for Neural Architecture Generation, called EDNAG, which simulates the denoising process in diffusion models within an evolutionary paradigm. EDNAG combines the superior generation quality of diffusion models with the conditional guidance capabilities of evolutionary algorithms.
  • We introduce a Fitness-guided Denoising (FD) strategy, which utilizes fitness instead of trained networks to progressively generate offspring architecture samples from previous ones, achieving the first training-free conditional neural architecture generation approach.
  • Extensive experiments have demonstrated the SOTA performance of EDNAG in architecture generation with significantly fewer computational resources, as well as its outstanding adaptability and transferability.
2 Related Work
2.1 Neural Architecture Generation
Neural Architecture Generation (NAG) has evolved from VAE-based to GAN-based methods and has recently adopted diffusion-based techniques. NAVIGATOR-D3 Hemmi et al. (2024) is a VAE-based NAG method that extracts the graph features of architectures by an encoder and then reconstruct them through a decoder to generate optimal neural architectures. Following the paradigm of GAN, GA-NAS Such et al. (2020) trains a discriminator to distinguish winning architectures and a generator to produce architectures based on the rewards provided by the discriminator, constituting an adversarial optimization problem. DiffusionNAG represents the groundbreaking application of diffusion models in NAG, producing high-performing architectures through a denoising process with a trained score network. Rather than continuous graph diffusion framework as in DiffusionNAG, DiNAS leverages a discrete graph diffusion model with a score network to generate neural architectures. However, both DiffusionNAG and DiNAS usionNAG and DiNAS require numerous denoising iterations to generate architecture samples, along with frequent network inferences during iterations, leading to inefficiency in the generation process. Their iterative denoising process also suffers from slow convergence and a tendency to stagnate in local optima. Moreover, the aforementioned methods require the training process of one or more networks for generation tasks, which consume significant computational resources, while our method, EDNAG, uniquely achieves a network-free and training-free generation approach.
2.2 Evolutionary Algorithms
Evolutionary algorithms, designed for optimization problems, serve as population-based heuristic paradigms simulating the natural evolution of species, which are widely appreciated for their global optimization capabilities. Recent studies have revealed potential connections between generative models and evolutionary algorithms, and have begun preliminary attempts to address challenging optimization problems with diffusion models. Zhang et al. (2024) mathematically demonstrates the equivalence between diffusion and evolution, conceptualizing the diffusion process as reversed evolution and the denoising process as evolution. Hartl et al. (2024) improves evolutionary algorithms by integrating diffusion models, enabling the retention of historical information across generations and significantly boosting convergence efficiency. Indeed, the aforementioned studies have largely focused on leveraging the properties of diffusion models to enhance evolutionary algorithms for optimization tasks. However, no prior work has attempted to integrate the principles of evolutionary algorithms into diffusion models for conditional generation tasks. Meanwhile, neural architecture generation (NAG), as a conditional generation task aimed at producing optimal neural architectures, naturally highlights the potential of integrating evolutionary algorithms into neural architecture generation, forming a foundational motivation for EDNAG.
3 Method
In order to improve the efficiency and quality of architecture generation, EDNAG employs evolutionary algorithms to simulate the denoising process in diffusion models. Specifically, EDNAG utilizes fitness to guide the denoising process, as detailed in Section 3.2. Notably, the fitness of architecture samples is estimated by a neural predictor, which is detailed in Section 3.3.
3.1 Preliminary
3.1.1 Diffusion Models
Denoising Diffusion Implicit Model (DDIM) Song et al. (2020) is a representative diffusion model. It involves a forward diffusion process, adding Gaussian noise to original data, and a reverse denoising process, where the model learns to generate original data by reversing the diffusion process.
In the forward process, we blend data with Gaussian noise, perturbing the original data $x_{0}$ to the final diffused data $x_{T}$ as follows:
$x_{t}=\sqrt{\overline{\alpha_{t}}}x_{0}+\sqrt{1-\overline{\alpha_{t}}}\epsilon_{t},\enspace\epsilon_{t}\sim\mathcal{N}(0,I),$ (1)
where $\overline{\alpha}_{t}=\prod_{i=1}^{T}\alpha_{i}$, representing the portion of data that remains undisturbed. In the reverse process, DDIM removes the noise from the randomly initialized data step by step, which is represented as:
$x_{t-1}=\sqrt{\overline{\alpha_{t-1}}}\underbrace{\left(\frac{x_{t}-\sqrt{1-\overline{\alpha_{t}}}\hat{\epsilon}_{\theta}\left(x_{t},t\right)}{\sqrt{\overline{\alpha_{t}}}}\right)}_{\text{predicted }x_{0}}+$ (2)
$\underbrace{\sqrt{1-\overline{\alpha_{t-1}}-\sigma_{t}^{2}}\cdot\hat{\epsilon}_{\theta}\left(x_{t},t\right)}_{\text{direction pointing to }x_{t}}+\underbrace{\sigma_{t}\epsilon_{t}}_{\text{random noise}}.$
where $\sigma_{t}$ controlls the degree of random perturbations $\epsilon_{t}\sim\mathcal{N}(0,I)$ during the denoising process. Notably, DDIM leverages a neural network $\epsilon_{\theta}$ to predict the perturbed noise $\hat{\epsilon}_{\theta}\left(x_{t},t\right)$ at time step $t$, guiding the denoising process. By iteratively repeating this denoising process in Equation 2, DDIM progressively generates the final samples. Our method, EDNAG, performs architecture generation based on the denoising process of DDIM, but eliminates the necessity for a network to estimate $\hat{\epsilon}_{\theta}\left(x_{t},t\right)$.
3.1.2 Evolutionary Algorithms
Evolutionary Algorithms (EAs) Coello Coello (2024); Song et al. (2023) are heuristic frameworks known for their superior optimization capabilities, where Genetic Algorithms (GAs) Alhijawi and Awajan (2024); Katoch et al. (2021) serve as a foundational and influential subset. GAs drive the evolution of populations through fitness-oriented natural selection. In each iteration, the generation of new individuals is facilitated by operators such as mutation and crossover, followed by a fitness-based selection process.
3.2 Evolutionary Diffusion-based NAG
3.2.1 Overview
EDNAG generates neural architectures through an iterative denoising process. In each iteration, EDNAG first evaluates the fitness of the architecture population using a neural predictor, then employs the fitness-guided denoising (FD) strategy, and finally leverages optimization strategies to select architecture samples. In the FD strategy, the estimation of the optimal architectures $\hat{x}_{0}$ is crucial to generate samples for the next generation. Specifically, the FD strategy maps samples $x_{t}$ to the probability density of $\hat{x}_{0}$ based on their fitness, thereby estimating $\hat{x}_{0}$ and then executing denoising iterations. The pipeline of EDNAG is described in  Figure 3, and the pseudocode is displayed in  Algorithm 1.
Figure 3: Fitness-guided denoising process. EDNAG generates neural architectures through iterative denoising process in FD strategy. In each iteration, FD maps the fitness of $x_{t}$ to the probability density of optimal samples $x_{0}$, subsequently uses the DDIM denoising formula to obtain $x_{t-1}$, and finally employs optimization strategies to select the final samples $x_{t-1}$ from $x_{t}$ and denoised $x_{t-1}$.
Algorithm 1 Neural Architecture Generation Process
Input: Population size $N$, number of nodes $D_{1}$, types of operations $D_{2}$, denoising steps $T$, fitness evaluator $f$, mapping function $g$.
Output: Generated optimal neural architectures.
1:Sample $x_{T}\sim\mathcal{N}(0,I^{N\times D_{1}\times D_{2}})$.
2:for $t$ in $[T,T-1,\ldots,1]$ do
3:     Map fitness $f(x_{t})$ to probability desity $g[f(x_{t})]$.
4:     Denoise from $x_{t}$ to $\hat{x}_{t-1}$ according to $g[f(x_{t})]$.
5:     Select final samples $x_{t-1}$ at the $t-1$-th time step from parent $x_{t}$ and denoised offspring $\hat{x}_{t-1}$ by optimization strategies.
6:end for
7:return $x_{0}$
In the following sections, we first explain the fitness-guided denoising (FD) strategy. Next, we describe our optimization strategies for selecting architecture samples.
3.2.2 Denoising Process of EDNAG
FD strategy simulates the denoising process of DDIM to derive new samples $x_{t-1}$ from previous samples $x_{t}$, generating offspring samples and driving the evolutionary process towards optimal solutions. By iteratively performing $T$ steps of denoising in FD strategy, the initial architecture samples $x_{T}$ (random noise samples) can be progressively transformed into optimal architecture samples $x_{0}$. For each denoising iteration, let $\hat{x}_{0}$ denotes the predicted $x_{0}$ in  Equation 2:
$\hat{x}_{0}=\frac{x_{t}-\sqrt{1-\overline{\alpha_{t}}}\hat{\epsilon}_{\theta}\left(x_{t},t\right)}{\sqrt{\overline{\alpha_{t}}}}.$ (3)
With Equation 3, the denoising process can be transformed as Equation 4.
$x_{t-1}$ $=\sqrt{\overline{\alpha_{t-1}}}\hat{x}_{0}+\sigma_{t}\epsilon_{t}$ (4)
$+$ $\sqrt{1-\overline{\alpha_{t-1}}-\sigma_{t}^{2}}\cdot\frac{x_{t}-\sqrt{\overline{\alpha_{t}}}\hat{x}_{0}}{\sqrt{1-\overline{\alpha_{t}}}}.$
Therefore, to perform the denoising process from $x_{t}$ to $x_{t-1}$, it is only necessary to determine the predicted optimal architectures $\hat{x}_{0}$ based on $x_{t}$, $\overline{\alpha_{t}}$ and $t$. Consequently, we propose a fitness-based FD strategy to estimate $\hat{x}_{0}$.
Before diving into the FD strategy, we compare the generation process of EDNAG with that of traditional diffusion-based methods. Traditional diffusion models in Equation 2, adopted by DiifusionNAG and DiNAS, utilizes $\hat{\epsilon}{\theta}\left(x{t},t\right)$—predicted by a trained network—to guide the generation of subsequent samples. In contrast, EDNAG’s generative process, formulated in Equation 4, is driven by $\hat{x}_{0}$, which is estimated through the training-free fitness-based denoising strategy proposed in this work.
To estimate $\hat{x}_{0}$, FD strategy maps the distribution of $x_{t}$ to the distribution of $\hat{x}_{0}$ with fitness guidance. Specifically, we view the denoising process of DDIM as the evolutionary process of GAs, both transforming samples from an initial random Gaussian distribution to an ultimate optimal distribution progressively Zhang et al. (2024). We consider that individuals with higher fitness are more likely to be retained during evolution, resulting in a higher probability of appearing in the final samples $\hat{x}_{0}$. Therefore, higher fitness corresponds to a higher probability density in $\hat{x}_{0}$. We model this relationship as a mapping function $g(x)$, which maps the fitness $f(x_{t})$ of $x_{t}$ to the probability density $p(\hat{x}_{0})$ of $\hat{x}_{0}$:
$p(\hat{x}_{0}=x)=g[f(x)].$ (5)
Hence, the optimal architectures $\hat{x}_{0}$ can be estimated from the samples $x_{t}$ at time step $t$ and their fitness $f(x_{t})$, as detailed below:
$p(\hat{x}_{0}=x|x_{t})=\frac{p(\hat{x}_{0}=x)\cdot p(x_{t}|x_{0}=x)}{p(x_{t})},$ (6)
where $p(x_{t}|x_{0}=x)$ is s derived from the diffusion process described by Equation 1, as shown below:
$p(x_{t}|x_{0}=x)\sim\mathcal{N}(x_{t};\sqrt{\overline{\alpha}_{t}}x,1-\overline{\alpha}_{t}).$ (7)
We further define $x_{t}$ as the previous samples consisting of $N$ architectures, where $x_{t}=[\boldsymbol{x}_{t}^{1},\boldsymbol{x}_{t}^{2},\ldots,\boldsymbol{x}_{t}^{N}]$. Combining Equations 6 and 7, $\hat{x}_{0}$ can be estimated as follows:
$\hat{x}_{0}=\frac{1}{p(x_{t})}\sum_{\boldsymbol{x}_{t}^{i}\in x_{t}}{g[f(\boldsymbol{x}_{t}^{i})]\mathcal{N}(x_{t};\sqrt{\overline{\alpha_{t}}}\boldsymbol{x}_{t}^{i},1-\overline{\alpha_{t}})\boldsymbol{x}_{t}^{i}},$ (8)
$p(x_{t})=\sum_{\boldsymbol{x}_{t}^{i}\in x_{t}}{g[f(\boldsymbol{x}_{t}^{i})]\mathcal{N}(x_{t};\sqrt{\overline{\alpha_{t}}}\boldsymbol{x}_{t}^{i},1-\overline{\alpha_{t}})},$
where $p(x_{t})$ actually serves as a regularization term. In fact, $\hat{x}_{0}$ can be regarded as the weighted summation of each neural architecture $\boldsymbol{x}_{t}^{i}$ in samples $x_{t}$, where architectures with higher fitness have a greater influence on $\hat{x}_{0}$.
Finally, with Equations 4 and LABEL:main_estimation_x_0, we can denoise from $x_{t}$ to $x_{t-1}$ as follows:
[…]
By iteratively applying the FD denoising strategy across $x_{t}$, $x_{t-1}$,…, $x_{0}$, which guides the evolution of architecture samples toward higher-fitness subspaces, EDNAG ultimately generates the global-optimal neural architectures without trained networks.
3.2.3 Optimization Strategies for Selection
In each denoising iteration, we apply LABEL:final_total_denoise to denoise architecture samples from $x_{t}$ to $x_{t-1}$. Subsequently, we employ optimization strategies (denoted as $S$) to select the final $x_{t-1}$ at time step $t-1$ from previous samples $x_{t}$ and denoised samples $x_{t-1}$:
$x_{t-1}=S\left(x_{t},x_{t-1}\right).$ (10)
Specifically, we adopt an elitist strategy Kaveh et al. (2022) to retain the fittest architectures, a diversity strategy to preserve diverse architectures Brázda et al. (2021), and finally the roulette wheel strategy Lipowski and Lipowska (2012) to select remaining architectures. These optimization strategies, which simulate natural selection in GAs, ensure task-optimality while simultaneously improving the stability of the generation process and promoting diversity within the generated architecture samples.
3.3 Transferable Neural Predictor for Guidance
Figure 4: Neural predictor in EDNAG. Our predictor comprises coreset selection, dataset encoding, architecture encoding, and accuracy prediction modules.
For task-specific conditional neural architecture generation, fitness evaluation is critical in guiding the denoising process of generation, which is conducted by a transferable neural predictor. Following DiffusionNAG, our neural predictor comprises three components: a set encoder $p(z_{\mathcal{D}}|\mathcal{D})$Lee et al. (2019), a graph encoder $p(z_{\mathcal{G}}|\mathcal{G})$Zhang et al. (2019), and a linear predictor $p(f(x)|z_{\mathcal{D}},z_{\mathcal{G}})$. In DiffusionNAG, the set encoder $p(z_{\mathcal{D}}|\mathcal{D})$ encodes a small subset of images randomly sampled from the task dataset for each prediction. This naive random sampling for each prediction overlooks the uniformity of dataset distribution, resulting in biased and unstable predictions. To address these issues, we first select a coreset Liu et al. (2023); Guo et al. (2022) from the task dataset and feed it into the set encoder $p(z_{\mathcal{D}}|\mathcal{D})$. Specifically, during prediction, each class of images in the task dataset $\mathcal{T}=\left\{\left(x_{t}^{i},y_{t}^{i}\right)\right\}{i=1}^{|\mathcal{T}|}$ undergoes K-means clustering to select centroids of clusters as coreset $\mathcal{D}=\left\{\left(x_{t}^{i},y_{t}^{i}\right)\right\}{i=1}^{|\mathcal{D}|}$, where $|\mathcal{D}|$ represents the number of samples per class and $|\mathcal{D}|\ll|\mathcal{T}|$. Subsequently, the coreset and the neural architecture are encoded by the set encoder $p(z_{\mathcal{D}}|\mathcal{D})$ and the graph encoder $p(z_{\mathcal{G}}|\mathcal{G})$ respectively to the latent representations $z_{\mathcal{D}}$ and $z_{\mathcal{G}}$. Finally, the linear predictor predicts the performance of the neural architecture. The entire process is detailed in Algorithm 2.
Algorithm 2 Performance Prediction for Neural Architectures
Input: Neural architecture $\mathcal{G}$, task dataset $\mathcal{T}$, number of samples per class $n$
Output: Estimated fitness $f(x)$ of architectures.
1:Apply K-means clustering to each class in original task dataset $\mathcal{T}$ to form $n$ clusters.
2:Select cluster centroids as core set $\mathcal{D}$, where $|\mathcal{D}|=n$.
3:for $\mathcal{G}$ in $[x^{1},x^{2},\ldots,x^{N}]$ do
4:     Encode core set $\mathcal{D}$ as $z_{\mathcal{D}}$ by $p(z_{\mathcal{D}}|\mathcal{D})$.
5:     Encode neural architecture $\mathcal{G}$ as $z_{\mathcal{G}}$ by $p(z_{\mathcal{G}}|\mathcal{G})$.
6:     Predict performance $f(x)$ by $p(f(x)|z_{\mathcal{D}},z_{\mathcal{G}})$.
7:end for
8:return list of estimated fitness $f(x)$
Notably, similar to previous methods, our neural predictor requires training. Although EDNAG shares this commonality with other NAG approaches based on generative models with proxy predictors, its key distinction and advantage lie in the generative model, which operates without score networks, thus significantly enhancing efficiency.
4 Experiments
In Section 4.1, we conduct sufficient experiments in NAS-BENCH-101 and NAS-BENCH-201 search spaces and extend our study to TransNASBench-101-Micro/Macro and MobileNetV3 search space. Subsequently, we perform ablation experiments on the FD strategy, as detailed in Section 4.2. Finally, further experiments on generation efficiency and neural predictors are carried out in Section 4.3.
4.1 Experiments on Neural Architecture Generation
4.1.1 Experiment Setup
We compare EDNAG with mainstream NAS methods, demonstrating its superior capabilities in architecture generation across various datasets. Classic NAS includes random search-based NAS, evolutionary-based NAS, and reinforcement learning-based NAS methods. One-shot NAS performs NAS within sub-networks of a supernet, including weight inheritance-based and gradient-based methods. BO-based NAS employs Bayesian optimization techniques (e.g. Gaussian processes) to search for optimal neural architectures. NAG employs different generative models to directly generate neural architectures.
For the generation process, the denoising step of EDNAG is 100. The noise scale of diffusion ranges from $\alpha_{T}=10^{-4}$ to $\alpha_{0}=1-10^{-4}$ and the noise scale of denoising is $\sigma=0.8$. The population size (or batch size) is 100 for MobileNetV3 search space and 30 for other search spaces.
4.1.2 Experiments on NAS-Bench-101 Search Space
Table 1: Experiments on NAS-BENCH-101 search space. Metrics are reported over 3 runs on CIFAR-10 with 95% confidence intervals.
Method Val Acc Test Acc
RS Bergstra and Bengio (2012) 93.92$\pm$0.11 93.42$\pm$0.02
RL Williams (1992) 94.23$\pm$0.33 93.72$\pm$0.27
RE Real et al. (2019) 94.28$\pm$0.24 93.78$\pm$0.16
FADZS-NAS Wang et al. (2024) 93.80$\pm$0.08 93.22$\pm$0.08
DiNAS Asthana et al. (2024) 94.83$\pm$0.03 94.40$\pm$0.07
EDNAG (Ours) 94.86$\pm$0.00 94.46$\pm$0.01
Global Best 94.86 94.47
NAS-Bench-101 is a typical search space encompassing $4.23\times 10^{5}$ unique neural architectures. It provides up to 7 layers with 3 candidate operations and a maximum of nine inter-layer connections.
We conduct experiments on NAS-Bench-101 for image classification task of CIFAR-10, displayed in Table 1. EDNAG outperforms previous methods, particularly surpassing the prior state-of-the-art approach (DiNAS), by improving validation and test accuracy by 0.03% and 0.06% respectively, which highlights the strong capability of EDNAG in generating task-optimal architectures.
4.1.3 Experiments on NAS-Bench-201 Search Space
Table 2: Experiments on NAS-BENCH-201 search space. Acc refers to the average accuracy over 3 runs with 95% confidence intervals, and Archs denotes the number of neural architectures actually trained to achieve this accuracy.
Type Method CIFAR-10 CIFAR-100 Aircraft Pets
Acc. (%) Archs Acc. (%) Archs Acc. (%) Archs Acc. (%) Archs
Classic NAS REINFORCE Williams (1992) 93.85$\pm$0.37 $>$ 500 71.71$\pm$1.09 $>$ 500 - - - -
RS Bergstra and Bengio (2012) 93.70$\pm$0.36 $>$ 500 71.04$\pm$1.07 $>$ 500 - - - -
REA Real et al. (2019) 93.92$\pm$0.30 $>$ 500 71.84$\pm$0.99 $>$ 500 - - - -
One-shot NAS RSPS Li and Talwalkar (2019) 84.07$\pm$3.61 N/A 52.31$\pm$5.77 N/A 42.19$\pm$3.88 N/A 22.91$\pm$1.65 N/A
SETN Dong and Yang (2019a) 87.64$\pm$0.00 N/A 59.09$\pm$0.24 N/A 44.84$\pm$3.96 N/A 25.17$\pm$1.68 N/A
GDAS Dong and Yang (2019b) 93.61$\pm$0.09 N/A 70.70$\pm$0.30 N/A 53.52$\pm$0.48 N/A 24.02$\pm$2.75 N/A
PC-DARTS Xu et al. (2020) 93.66$\pm$0.17 N/A 66.64$\pm$2.34 N/A 26.33$\pm$3.40 N/A 25.31$\pm$1.38 N/A
DrNAS Chen et al. (2021) 94.36$\pm$0.00 N/A 73.51$\pm$0.00 N/A 46.08$\pm$7.00 N/A 26.73$\pm$2.61 N/A
BO-based NAS BOHB Falkner et al. (2018) 93.61$\pm$0.52 $>$ 500 70.85$\pm$1.28 $>$ 500 - - - -
BANANAS White et al. (2021) 94.37$\pm$0.00 46 73.51$\pm$0.00 88 41.72$\pm$0.00 40 40.15$\pm$1.59 17
NASBOWL Ru et al. (2021) 94.34$\pm$0.00 100 73.51$\pm$0.00 87 53.73$\pm$0.83 40 41.29$\pm$1.10 17
HEBO Cowen-Rivers et al. (2022) 94.34$\pm$0.00 100 72.62$\pm$0.20 100 49.32$\pm$6.10 40 40.55$\pm$1.15 18
TNAS Shala et al. (2023) 94.37$\pm$0.00 29 73.51$\pm$0.00 59 59.15$\pm$0.58 26 40.00$\pm$0.00 6
NAG DiNAS Asthana et al. (2024) 94.37$\pm$0.00 192 73.51$\pm$0.00 192 - - - -
MetaD2A Lee et al. (2021) 94.37$\pm$0.00 100 73.34$\pm$0.04 100 57.71$\pm$0.20 40 39.04$\pm$0.20 40
DiffusionNAG An et al. (2023) 94.37$\pm$0.00 1 73.51$\pm$0.00 2 58.83$\pm$3.75 3 41.80$\pm$3.82 2
EDNAG (Ours) 94.37$\pm$0.00 1 73.51$\pm$0.00 1 60.14$\pm$0.89 1 46.17$\pm$0.67 1
Global Best 94.37 N/A 73.51 N/A - N/A - N/A
NAS-BENCH-201 is a widely used cell-based search space comprising 15,625 distinct neural architectures. Each cell in NAS-Bench-201 includes 6 layers and 5 predefined operations (zeroize, skip-connect, 1×1 conv, 3×3 conv, 3×3 avg pool) for each layer.
We perform conditional generation of architectures across four diverse task datasets within the NAS-BENCH-201 search space, as shown in Table 2. EDNAG generates neural architectures and subsequently trains the top-performing architecture on the task datasets to validate their performance. In experiments on CIFAR-10 and CIFAR-100, EDNAG achieves comparable performance on test accuracies compared to previous baselines. On Aircraft and Pets, EDNAG surpasses previous baselines with improvements of 2.23% and 10.45% on test accuracies, respectively. Across diverse task datasets, EDNAG outperforms all previous NAS and NAG methods, achieving state-of-the-art (SOTA) performance and showcasing exceptional capability in generating optimal architectures. This is attributed to the FD strategy, which efficiently guides architecture samples toward high-fitness subspaces during the iterative generation process, serving as a powerful conditional mechanism to steer architectures toward global optimality. Compared to DiffusionNAG, the previous SOTA baseline that directly generates neural architectures using naive diffusion models with limited optimization abilities, EDNAG excels with its superior global optimization capabilities while maintaining efficient generative abilities.
Moreover, EDNAG also significantly improves the efficiency of architecture validation by reducing the number of architectures (Archs) requiring training to achieve final accuracy. For CIFAR-10 and CIFAR-100, the final accuracies are obtained from the NAS-Bench-201 benchmarks Dong and Yang (2020) by querying only the top-1 generated architecture. For Aircraft and Pets, only the top-1 architecture requires full training in EDNAG, enhancing validation efficiency by 3$\times$ and 2$\times$ compared to DiffusionNAG. Results illustrate that the top-1 neural architecture identified by our neural predictor attains optimal accuracy after training and testing, demonstrating its remarkable performance, achieved by the application of coreset selection prior to dataset encoding. Unlike the random sampling method used in DiffusionNAG, this approach significantly enhances the stability and reliability of the predicted accuracies.
4.1.4 Experiments on TransNASBench-101 Search Space
Table 3: Experiments on TransNASBench-101 Search Space across seven tasks. Metrics are reported with 95% confidence intervals over 3 runs. For improved readability, the values in the table for SSIM, mIoU, and L2 Loss are scaled by 100.
Level Method Cls. Object Cls. Scene Autoencoding Surf. Normal Sem. Segment. Room Layout Jigsaw
Acc. (%) Acc. (%) SSIM SSIM mIoU L2 Loss Acc. (%)
Micro RS Bergstra and Bengio (2012) 45.16$\pm$0.4 54.41$\pm$0.3 55.94$\pm$0.8 56.85$\pm$0.6 25.21$\pm$0.4 61.48$\pm$0.8 94.47$\pm$0.3
REA Real et al. (2019) 45.39$\pm$0.2 54.62$\pm$0.2 56.96$\pm$0.1 57.22$\pm$0.3 25.52$\pm$0.3 61.75$\pm$0.8 94.62$\pm$0.3
PPO Schulman et al. (2017) 45.19$\pm$0.3 54.37$\pm$0.2 55.83$\pm$0.7 56.90$\pm$0.6 25.24$\pm$0.3 61.38$\pm$0.7 94.46$\pm$0.3
CATCH Chen et al. (2020) 45.27$\pm$0.5 54.38$\pm$0.2 56.13$\pm$0.7 56.99$\pm$0.6 25.38$\pm$0.4 60.70$\pm$0.7 -
EDNAG (Ours) 46.32$\pm$0.0 54.94$\pm$0.0 57.72$\pm$0.0 59.62$\pm$0.0 26.27$\pm$0.0 58.71$\pm$0.0 95.37$\pm$0.0
Global Best 46.32 54.94 57.72 59.62 26.27 58.71 95.37
Macro RS Bergstra and Bengio (2012) 46.85$\pm$0.3 56.50$\pm$0.4 70.06$\pm$3.1 60.70$\pm$0.9 28.37$\pm$0.5 59.35$\pm$1.0 96.78$\pm$0.2
REA Real et al. (2019) 47.09$\pm$0.4 56.57$\pm$0.4 69.98$\pm$3.6 60.88$\pm$1.0 28.87$\pm$0.4 58.73$\pm$1.1 96.88$\pm$0.2
PPO Schulman et al. (2017) 46.84$\pm$0.4 56.48$\pm$0.3 70.92$\pm$3.2 60.82$\pm$0.8 28.31$\pm$0.5 58.84$\pm$1.1 96.76$\pm$0.2
CATCH Chen et al. (2020) 47.29$\pm$0.3 56.49$\pm$0.3 70.36$\pm$3.0 60.85$\pm$0.7 28.71$\pm$0.4 59.37$\pm$0.6 -
EDNAG (Ours) 47.96$\pm$0.0 57.48$\pm$0.0 76.88$\pm$0.0 64.35$\pm$0.0 29.66$\pm$0.0 55.68$\pm$0.0 97.02$\pm$0.0
Global Best 47.96 57.48 76.88 64.35 29.66 55.68 97.02
TransNASBench-101 extends downstream tasks beyond basic image classification, encompassing a diverse set of 7 tasks: Object Classification, Scene Classification, Autoencoding, Surface Normal Estimation, Semantic Segmentation, Room Layout Prediction, and Jigsaw Puzzle Solving. Moreover, it defines two distinct types of search spaces: a cell-based micro search space and a stack-based macro skeleton search space, which encompass 4,096 and 3,256 unique architectures, respectively.
We conduct experiments in both micro-level and macro-level search spaces, as presented in Table 3. Across both search spaces, EDNAG exhibits significant performance improvements compared to previous baselines and consistently achieves globally optimal results across all downstream tasks while maintaining remarkable stability across three independent runs. These findings not only demonstrate the effectiveness of the FD strategy in improving global optimization capabilities for generative models, but also highlight the transferability of EDNAG for cross-task neural architecture generation within various downstream scenarios.
4.1.5 Experiments on MobileNetV3 Search Space
Table 4: Experiments on the MobileNetV3 search space. We report the average accuracy of different methods over 3 runs with with 95% confidence intervals.
Methods CIFAR-10 CIFAR-100 Aircraft Pets
MetaD2A 97.28$\pm$0.01 85.56$\pm$0.02 81.19$\pm$0.11 94.55$\pm$0.03
TNAS 97.22$\pm$0.00 85.30$\pm$0.04 80.86$\pm$0.15 94.47$\pm$0.10
DiffusionNAG 97.39$\pm$0.01 85.74$\pm$0.04 81.47$\pm$0.05 94.75$\pm$0.10
EDNAG (Ours) 97.40$\pm$0.01 85.81$\pm$0.02 81.87$\pm$0.10 95.06$\pm$0.14
Due to the relatively limited scale of the above search spaces, our study further extends to the larger and more complex MobileNetV3 search space to demonstrate the scalability of EDNAG. MobileNetV3 is a large search space which encompasses over $2.17\times 10^{19}$ distinct neural architectures, making it more than $10^{15}$ times larger than the NAS-Bench-201 search space. MobileNetV3 offers five stages, each consisting of 2 to 4 building blocks, where each building block is defined by the kernel size of convolutional layers chosen from {3, 5, 7} and the expansion rate of output channels chosen from {3, 4, 6}.
We conduct experiments on four distinct task datasets within the MobileNetV3 search space, as shown in Table 4. In CIFAR-10, CIFAR-100, Aircraft, and Pets, EDNAG demonstrates an improvement in accuracy of up to 0.49%. This experiment not only demonstrates the superior ability of architecture generation for EDNAG but also confirms its adaptability and effectiveness in large-scale search spaces. This is attributed to the fact that when transitioning to different search spaces, DiffusionNAG necessitates retraining the score network within the diffusion model to accommodate the noise matrix of the new dimensionality, which is troublesome and time-consuming. In contrast, as a training-free NAG method, EDNAG avoids the need for training or retraining, thereby enabling adaptation to new search spaces with negligible costs.
4.2 Ablation Study
Figure 5: Comparison of architecture generation using naive diffusion models, diffusion models with FD strategy, and diffusion models with both FD and optimization strategies.
The Fitness-guided Denoising (FD) strategy introduces evolutionary algorithms to emulate the iterative denoising process in diffusion models, which constitutes the core mechanism of EDNAG. Furthermore, the optimization strategy, driven by selection operators, further enhances the ability of EDNAG to generate globally optimal neural architectures. To deeply analyze their critical roles, we conducted a comparative analysis of the architecture generation process under three distinct conditions: naive diffusion models, diffusion models with the FD strategy, and diffusion models with the FD strategy and optimization strategy, as illustrated in Figure 5.
Naive diffusion models often results in convergence to local optima, resulting in suboptimal architecture generation. The integration of the FD strategy significantly enhances the accurancy of generated architectures, facilitating the generation of globally optimal solutions. In addition, the incorporation of the optimization strategy substantially accelerates the convergence of the denoising process. These results underscore the effectiveness of both the FD strategy and the optimization strategy within EDNAG, demonstrating their pivotal roles in achieving task-optimal architecture generation.
4.3 In-Depth Analysis
4.3.1 Comparison of Generation Efficiency
Table 5: Comparison of GPU time costs. We report the training time (hours) and inference time (seconds) for the architecture generation process.
Evaluation Metric EDNAG DiffusionNAG
Training Time (h) 0 3.43
Generation Time (s) CIFAR-10 2.5 257
CIFAR-100 2.5 261
Aircraft 57 274
Pets 38 263
As the first network-free neural architecture generation method, EDNAG boasts remarkable characteristics, including minimal computational resource consumption and rapid conditional generation, etc. To demonstrate its efficiency, we compare EDNAG with DiffusionNAG in terms of computational costs during both training and inference processes in the NAS-Bench-201 search space, as shown in Table 5. Since EDNAG is free of training, it takes 0 GPU hours for the training process, while DiffusionNAG requires 3.43 GPU hours in the NAS-BENCH-201 search space on a Tesla V100-SXM2. For generation process, we evaluate the inference time of both EDNAG and DiffusionNAG on an NVIDIA RTX 4090. EDNAG achieves speedups of up to 100$\times$, 100$\times$, 7$\times$, and 5$\times$ compared to DiffusionNAG on CIFAR-10, CIFAR-100, Aircraft and Pets, respectively. Therefore, EDNAG not only enables neural architecture generation without time-consuming training processes but also significantly reduces computational cost during the inference process. This is attributed to the FD strategy during the generative process, which directs the denoising process without relying on a score network, thereby achieving remarkable execution efficiency. In contrast, DiffusionNAG depends on a complex score network at each step of the generative process to predict the noise to be removed, which demands extensive and frequent network inference, significantly slowing down the generation process.
4.3.2 Quality Analysis of Generated Architectures
Figure 6: Distributions of generated architectures. In the NAS-BENCH-201 search space and CIFAR-10 dataset, we generate 1000 neural architectures using three different methods and visualize their accuracy distributions.
We generate 1000 neural architectures by EDNAG and DiffusionNAG, then we display their distribution in Figure 6. Compared to DiffusionNAG, EDNAG demonstrates a higher proportion of high-performing architectures. This demonstrates that the architectures generated by EDNAG exhibit a denser distribution within high-accuracy subspaces, reflecting a marked improvement in the quality of architecture generation. This advancement is achieved through the powerful conditional guidance and superior optimization ability by combining evolutionary algorithms, which contrasts sharply with the straightforward use of naive diffusion models in DiffusionNAG.
4.3.3 Analysis of Neural Predictor
Figure 7: Performance analysis of neural predictor. We predict the accuracy of architectures for CIFAR-10 and CIFAR-100, and display the Pearson correlation coefficient between predicted and actual accuracies.
Figure 8: Stability analysis of neural predictor. With neural predictors in DiffusionNAG and EDNAG, we assessed the accuracy of identical neural architectures on consistent task datasets across 100 repeated runs.
We conduct experiments within the NAS-BENCH-201 search space and report Pearson correlation coefficients between predicted and actual accuracies, as illustrated in Figure 7. On CIFAR-10, the predictor of DiffusionNAG suffers from biased predictions due to random sampling, resulting in a lower correlation coefficient, while EDNAG achieves a 12.09% improvement. On CIFAR-100, the predictor of EDNAG achieves a 7.29% enhancement compared to DiffusionNAG. This experiment reveals that the neural predictor of EDNAG, which utilizes K-Means clustering to select coreset images for dataset encoding, aligns more closely with real architectural accuracy. This approach effectively circumvents the limitations of random sampling in DiffusionNAG, which fails to preserve the uniformity of the target dataset distribution, resulting in biased predictions.
Moreover, during the denoising iterations, identical architectures in successive generations ($x_{t}$ and $x_{t-1}$) necessitate multiple evaluations of the same architecture. Our experiments, shown in Figure 8, evaluate the same neural architecture over 100 repeated runs. The results demonstrate significant fluctuations in the predicted accuracy of DiffusionNAG due to random image sampling. In contrast, EDNAG demonstrates consistent predicted accuracy, effectively avoiding random perturbations in fitness values and further enhancing the stability of the whole generation process.
5 Conclusion
This study introduces an Evolutionary Diffusion-based Neural Architecture Generation Approach, called EDNAG, an efficient neural architecture generation method without networks. We propose a fitness-guided denoising strategy (FD) that maps fitness to probabilistic space, thereby guiding the DDIM-based denoising process for architecture generation. Extensive experiments demonstrate that EDNAG achieves SOTA performance in generating task-optimal architectures and surpasses existing baselines in generation efficiency and computational cost. We believe that EDNAG will significantly contribute to the development of NAS by reducing computational costs and accelerate its practical application.

### v3

@@
 Furthermore, it eliminates the need for time-consuming training and boosts inference speed by an average of 50×, showcasing its exceptional efficiency.
+Codes are available at https://github.com/cloud-zhoubingye/EDNAG.git.
 1 Introduction
@@
 (2024); Lukasik et al.
-(2022) has revolutionized this optimization challenge by re-framing it from an optimization task into a generative task, thus mitigating the limitations imposed by large search spaces.
-The NAG framework leverages advanced generation techniques to directly produce optimized neural architectures that are finely tuned for specific downstream applications, offering a more scalable and resource-efficient solution.
+(2022) has revolutionized this optimization challenge by reframing it from an optimization task into a generation task, thus mitigating the limitations imposed by large search spaces.
+The NAG framework leverages advanced generation techniques to directly produce optimized neural architectures that are finely tuned for specific downstream applications, offering more scalable and resource-efficient solutions.
 Previous architecture generation techniques can be categorized into two primary classes: Variational Auto-Encoders (VAE)-based Lin et al.
@@
 (2023) methods.
-VAE-based methods train an encoder to map architecture features into latent space and a decoder to generate them from latent space, while GAN-based methods employ a generator to produce architectures with the reward from a discriminator.
+VAE-based methods train an encoder to map architecture features into embedding representations and a decoder to generate them from the latent space, while GAN-based methods employ a generator to produce architectures with the reward from a discriminator.
 However, both VAE-based and GAN-based methods require the simultaneous training of two networks, leading to significant challenges in high computational complexity, training stability Brock (2018), and generation quality.
@@
 (2024) employ a graph diffusion framework to generate architectures through iterative denoising steps, where each denoising iteration requires network inference to predict the noise to be eliminated.
-However, the generation process is inefficient due to the frequent network inferences and the multi-step denoising process, significantly slowing down the overall generation process.
-Moreover, when confronted with intricate and expansive search spaces, diffusion models often exhibit limitations in generating neural architectures with global optimality, as shown in Figure 2.
+However, the generation process is inefficient due to the frequent network inferences and the multi-step denoising process, which typically requires thousands of steps, significantly slowing down the generation process.
+Moreover, when confronted with complex and expansive search spaces, diffusion models often exhibit limitations in generating local optimal neural architectures.
 Their naive diffusion models frequently converge to local optima, leading to the generation of suboptimal architectures.
-Therefore, diffusion-based methods still encounter substantial computational complexity during the training phase and exhibit suboptimality in the generation of neural architectures.
-Figure 2: Each arrow in the figure represents a denoising step in diffusion models.
-DiffusionNAG simply adopts diffusion models, which tend to generate suboptimal architectures.
-In contrast, EDNAG leverages the global optimization capability of evolutionary algorithms to improve the generation process, making it more likely to produce globally optimal architectures.
+Therefore, diffusion-based methods still encounter substantial computational complexity during the training phase and exhibit sub-optimality in the generation of neural architectures.
 To the end, we propose a novel and efficient Evolutionary Diffusion-based Neural Architecture Generation Framework, called EDNAG.
@@
 Specifically, we introduce evolutionary algorithms to simulate the denoising process in diffusion models.
-Furthermore, we propose a Fitness-guided Denoising (FD) strategy to produces a new generation of architectures from the previous ones in each denoising iteration.
+Furthermore, we propose a Fitness-guided Denoising (FD) strategy to produce a new generation of architectures from the previous ones in each denoising iteration.
 Each newly produced architecture can be conceptualized as a weighted summation of prior samples, with superior architectures assigned higher weights to exert a more prominent influence.
 Therefore, architecture samples will progressively evolve towards high-fitness subspaces through multiple denoising iterations with FD strategy, thereby generating high-performing and global-optimal neural architectures.
-In particular, EDNAG performs task-specific neural architecture generation with a neural predictor, serving as a dataset-aware fitness evaluator, to guide the conditional generation process with the FD strategy.
+In particular, EDNAG performs task-specific neural architecture generation with a neural predictor, serving as a dataset-aware fitness evaluator to guide the conditional generation process within the FD strategy.
 Notably, the FD strategy eliminates the dependency on the score network in traditional diffusion models during the denoising process and achieves network-free neural architecture generation, therefore greatly improving generation efficiency.
-We performed extensive experiments within NAS-Bench-101 Ying et al.
-(2019) on CIFAR10 and NAS-Bench-201 Dong and Yang (2020) across four task datasets.
-These experiments demonstrate that EDNAG not only achieves state-of-the-art (SOTA) performance in architecture generation with up to 10.45% improvement in accuracy, but also exhibits an average 50× speedup compared to previous baselines.
-Furthermore, we expanded the search space to TransNASBench-101 Duan et al.
+We perform extensive experiments within NAS-Bench-101 Ying et al.
+(2019), NAS-Bench-201 Dong and Yang (2020), and NAS-Bench-301 Zela et al.
+(2020), demonstrating that EDNAG not only achieves state-of-the-art (SOTA) performance in architecture generation with up to 10.45% improvement in accuracy, but also exhibits an average 50$\times$ speedup compared to previous baselines.
+Furthermore, we expand the search space to TransNASBench-101 Duan et al.
 (2021) and MobileNetV3 Howard et al.
 (2019).
-The former search space involves seven distinct downstream computer vision tasks, while the latter features an extremely large search space.
-Experiments demonstrate that EDNAG consistently achieves state-of-the-art performance in architecture generation, highlighting its adaptability to larger search spaces and its transferability to diverse downstream tasks.
-Our code is available at https://github.com/cloud-zhoubingye/EvoDiff.git.
-In summary, our contributions are as follows:
+The former involves seven distinct downstream computer vision tasks, while the latter features an extremely large search space.
+Experiments demonstrate EDNAG’s adaptability to expansive search spaces and its transferability to diverse downstream tasks.
+In summary, our contributions are as follows.
   • We propose a novel Evolutionary Diffusion-based framework for Neural Architecture Generation, called EDNAG, which simulates the denoising process in diffusion models within an evolutionary paradigm.
-EDNAG combines the superior generation quality of diffusion models with the conditional guidance capabilities of evolutionary algorithms.
-  • We introduce a Fitness-guided Denoising (FD) strategy, which utilizes fitness instead of trained networks to progressively generate offspring architecture samples from previous ones, achieving the first training-free conditional neural architecture generation approach.
+EDNAG combines the superior generation quality of diffusion models with the global optimization capabilities of evolutionary algorithms.
+  • We introduce a Fitness-guided Denoising (FD) strategy, which utilizes fitness instead of trained networks to progressively generate offspring architecture samples from previous ones, achieving the first network-free neural architecture generation approach.
   • Extensive experiments have demonstrated the SOTA performance of EDNAG in architecture generation with significantly fewer computational resources, as well as its outstanding adaptability and transferability.
@@
 NAVIGATOR-D3 Hemmi et al.
-(2024) is a VAE-based NAG method that extracts the graph features of architectures by an encoder and then reconstruct them through a decoder to generate optimal neural architectures.
+(2024) is a VAE-based NAG method that extracts the graph features of architectures by an encoder and then reconstructs them through a decoder to generate optimal neural architectures.
 Following the paradigm of GAN, GA-NAS Such et al.
@@
 DiffusionNAG represents the groundbreaking application of diffusion models in NAG, producing high-performing architectures through a denoising process with a trained score network.
-Rather than continuous graph diffusion framework as in DiffusionNAG, DiNAS leverages a discrete graph diffusion model with a score network to generate neural architectures.
-However, both DiffusionNAG and DiNAS usionNAG and DiNAS require numerous denoising iterations to generate architecture samples, along with frequent network inferences during iterations, leading to inefficiency in the generation process.
+Rather than a continuous graph diffusion framework as in DiffusionNAG, DiNAS leverages a discrete graph diffusion model with a score network to generate neural architectures.
+However, both DiffusionNAG and DiNAS require numerous denoising iterations to generate architecture samples, along with frequent network inferences during iterations, leading to inefficiency in the generation process.
 Their iterative denoising process also suffers from slow convergence and a tendency to stagnate in local optima.
@@
 Hartl et al.
-(2024) improves evolutionary algorithms by integrating diffusion models, enabling the retention of historical information across generations and significantly boosting convergence efficiency.
+(2024) improves evolutionary algorithms by integrating diffusion models, enabling the retention of historical information across generations.
 Indeed, the aforementioned studies have largely focused on leveraging the properties of diffusion models to enhance evolutionary algorithms for optimization tasks.
@@
 (2020) is a representative diffusion model.
-It involves a forward diffusion process, adding Gaussian noise to original data, and a reverse denoising process, where the model learns to generate original data by reversing the diffusion process.
+It involves a forward diffusion process, adding Gaussian noise to the original data, and a reverse denoising process, where the model learns to generate the original data by reversing the diffusion process.
 In the forward process, we blend data with Gaussian noise, perturbing the original data $x_{0}$ to the final diffused data $x_{T}$ as follows:
@@
 $\underbrace{\sqrt{1-\overline{\alpha_{t-1}}-\sigma_{t}^{2}}\cdot\hat{\epsilon}_{\theta}\left(x_{t},t\right)}_{\text{direction pointing to }x_{t}}+\underbrace{\sigma_{t}\epsilon_{t}}_{\text{random noise}}.$
-where $\sigma_{t}$ controlls the degree of random perturbations $\epsilon_{t}\sim\mathcal{N}(0,I)$ during the denoising process.
-Notably, DDIM leverages a neural network $\epsilon_{\theta}$ to predict the perturbed noise $\hat{\epsilon}_{\theta}\left(x_{t},t\right)$ at time step $t$, guiding the denoising process.
+where $\sigma_{t}$ controls the degree of random perturbations $\epsilon_{t}\sim\mathcal{N}(0,I)$ during the denoising process.
+Notably, DDIM leverages a neural network $\epsilon_{\theta}$ to predict perturbed noise $\hat{\epsilon}_{\theta}\left(x_{t},t\right)$ at time step $t$, guiding the denoising process.
 By iteratively repeating this denoising process in Equation 2, DDIM progressively generates the final samples.
-Our method, EDNAG, performs architecture generation based on the denoising process of DDIM, but eliminates the necessity for a network to estimate $\hat{\epsilon}_{\theta}\left(x_{t},t\right)$.
+Our method, EDNAG, performs architecture generation based on the denoising process of DDIM, but eliminates the need for a network to estimate $\hat{\epsilon}_{\theta}\left(x_{t},t\right)$.
 3.1.2 Evolutionary Algorithms
-Evolutionary Algorithms (EAs) Coello Coello (2024); Song et al.
-(2023) are heuristic frameworks known for their superior optimization capabilities, where Genetic Algorithms (GAs) Alhijawi and Awajan (2024); Katoch et al.
-(2021) serve as a foundational and influential subset.
+Evolutionary Algorithms (EAs) Coello Coello (2024) are heuristic frameworks known for their superior optimization capabilities, where Genetic Algorithms (GAs) Alhijawi and Awajan (2024) serve as a foundational and influential subset.
 GAs drive the evolution of populations through fitness-oriented natural selection.
@@
 Figure 2: Fitness-guided denoising process.
-EDNAG generates neural architectures through iterative denoising process in FD strategy.
-In each iteration, FD maps the fitness of $x_{t}$ to the probability density of optimal samples $x_{0}$, subsequently uses the DDIM denoising formula to obtain $x_{t-1}$, and finally employs optimization strategies to select the final samples $x_{t-1}$ from $x_{t}$ and denoised $x_{t-1}$.
+EDNAG generates neural architectures through iterative denoising process with FD strategy.
+In each iteration, FD strategy maps the fitness of $x_{t}$ to the probability density of optimal samples $x_{0}$, subsequently uses the DDIM denoising formula to obtain $x_{t-1}$, and finally employs optimization strategies to select the final samples $x_{t-1}$ from $x_{t}$ and denoised $x_{t-1}$.
 Algorithm 1 Neural Architecture Generation Process
@@
 2:for $t$ in $[T,T-1,\ldots,1]$ do
-3:     Map fitness $f(x_{t})$ to probability desity $g[f(x_{t})]$.
+3:     Map fitness $f(x_{t})$ to probability density $g[f(x_{t})]$.
 4:     Denoise from $x_{t}$ to $\hat{x}_{t-1}$ according to $g[f(x_{t})]$.
@@
 7:return $x_{0}$
-In the following sections, we first explain the fitness-guided denoising (FD) strategy. Next, we describe our optimization strategies for selecting architecture samples.
+In the following sections, we first explain the fitness-guided denoising (FD) strategy.
+Next, we describe our optimization strategies for selecting architecture samples.
+More detailed derivations and explanations are provided in the Appendix.
 3.2.2 Denoising Process of EDNAG
@@
 $\hat{x}_{0}=\frac{x_{t}-\sqrt{1-\overline{\alpha_{t}}}\hat{\epsilon}_{\theta}\left(x_{t},t\right)}{\sqrt{\overline{\alpha_{t}}}}.$ (3)
-With Equation 3, the denoising process can be transformed as Equation 4.
+With Equation 3, the denoising process can be transformed into Equation 4.
 $x_{t-1}$ $=\sqrt{\overline{\alpha_{t-1}}}\hat{x}_{0}+\sigma_{t}\epsilon_{t}$ (4)
@@
 Consequently, we propose a fitness-based FD strategy to estimate $\hat{x}_{0}$.
-Before diving into the FD strategy, we compare the generation process of EDNAG with that of traditional diffusion-based methods.
-Traditional diffusion models in Equation 2, adopted by DiifusionNAG and DiNAS, utilizes $\hat{\epsilon}{\theta}\left(x{t},t\right)$—predicted by a trained network—to guide the generation of subsequent samples.
-In contrast, EDNAG’s generative process, formulated in Equation 4, is driven by $\hat{x}_{0}$, which is estimated through the training-free fitness-based denoising strategy proposed in this work.
 To estimate $\hat{x}_{0}$, FD strategy maps the distribution of $x_{t}$ to the distribution of $\hat{x}_{0}$ with fitness guidance.
@@
 […]
-By iteratively applying the FD denoising strategy across $x_{t}$, $x_{t-1}$,…, $x_{0}$, which guides the evolution of architecture samples toward higher-fitness subspaces, EDNAG ultimately generates the global-optimal neural architectures without trained networks.
+By iteratively applying the FD denoising strategy across $x_{t}$, $x_{t-1}$,…, $x_{0}$, which guides the evolution of architecture samples toward higher-fitness subspaces, EDNAG ultimately generates global-optimal neural architectures without trained networks.
 3.2.3 Optimization Strategies for Selection
@@
 To address these issues, we first select a coreset Liu et al.
-(2023); Guo et al.
 (2023) from the task dataset and feed it into the set encoder $p(z_{\mathcal{D}}|\mathcal{D})$.
 Specifically, during prediction, each class of images in the task dataset $\mathcal{T}=\left\{\left(x_{t}^{i},y_{t}^{i}\right)\right\}{i=1}^{|\mathcal{T}|}$ undergoes K-means clustering to select centroids of clusters as coreset $\mathcal{D}=\left\{\left(x_{t}^{i},y_{t}^{i}\right)\right\}{i=1}^{|\mathcal{D}|}$, where $|\mathcal{D}|$ represents the number of samples per class and $|\mathcal{D}|\ll|\mathcal{T}|$.
-Subsequently, the coreset and the neural architecture are encoded by the set encoder $p(z_{\mathcal{D}}|\mathcal{D})$ and the graph encoder $p(z_{\mathcal{G}}|\mathcal{G})$ respectively to the latent representations $z_{\mathcal{D}}$ and $z_{\mathcal{G}}$.
+Subsequently, the coreset and the neural architecture are encoded by the set encoder $p(z_{\mathcal{D}}|\mathcal{D})$ and the graph encoder $p(z_{\mathcal{G}}|\mathcal{G})$, respectively, into latent representations $z_{\mathcal{D}}$ and $z_{\mathcal{G}}$.
 Finally, the linear predictor predicts the performance of the neural architecture.
@@
 4 Experiments
-In Section 4.1, we conduct sufficient experiments in NAS-BENCH-101 and NAS-BENCH-201 search spaces and extend our study to TransNASBench-101-Micro/Macro and MobileNetV3 search space.
+In Section 4.1, we conduct sufficient experiments in NAS-BENCH-101, NAS-BENCH-201, and NAS-Bench-301 search spaces and extend our study to TransNASBench-101-Micro/Macro and MobileNetV3 search spaces.
 Subsequently, we perform ablation experiments on the FD strategy, as detailed in Section 4.2.
 Finally, further experiments on generation efficiency and neural predictors are carried out in Section 4.3.
+For sensitivity analysis of parameters, comprehensive comparisons of GPU time across 14 baselines, and more experiments, please refer to the Appendix.
 4.1 Experiments on Neural Architecture Generation
@@
 The population size (or batch size) is 100 for MobileNetV3 search space and 30 for other search spaces.
-4.1.2 Experiments on NAS-Bench-101 Search Space
-Table 1: Experiments on NAS-BENCH-101 search space. Metrics are reported over 3 runs on CIFAR-10 with 95% confidence intervals.
-Method Val Acc Test Acc
-RS Bergstra and Bengio (2012) 93.92$\pm$0.11 93.42$\pm$0.02
-RL Williams (1992) 94.23$\pm$0.33 93.72$\pm$0.27
-RE Real et al. (2019) 94.28$\pm$0.24 93.78$\pm$0.16
-FADZS-NAS Wang et al. (2024) 93.80$\pm$0.08 93.22$\pm$0.08
-DiNAS Asthana et al. (2024) 94.83$\pm$0.03 94.40$\pm$0.07
-EDNAG (Ours) 94.86$\pm$0.00 94.46$\pm$0.01
-Global Best 94.86 94.47
-NAS-Bench-101 is a typical search space encompassing $4.23\times 10^{5}$ unique neural architectures.
-It provides up to 7 layers with 3 candidate operations and a maximum of nine inter-layer connections.
-We conduct experiments on NAS-Bench-101 for image classification task of CIFAR-10, displayed in Table 1.
-EDNAG outperforms previous methods, particularly surpassing the prior state-of-the-art approach (DiNAS), by improving validation and test accuracy by 0.03% and 0.06% respectively, which highlights the strong capability of EDNAG in generating task-optimal architectures.
+More implementation details, search spaces description, and introduction to baselines can be found in the Appendix.
 4.1.2 Experiments on NAS-Bench-201 Search Space
@@
 NAS-BENCH-201 is a widely used cell-based search space comprising 15,625 distinct neural architectures.
-Each cell in NAS-Bench-201 includes 6 layers and 5 predefined operations (zeroize, skip-connect, 1×1 conv, 3×3 conv, 3×3 avg pool) for each layer.
-We perform conditional generation of architectures across four diverse task datasets within the NAS-BENCH-201 search space, as shown in Table 2.
-EDNAG generates neural architectures and subsequently trains the top-performing architecture on the task datasets to validate their performance.
-In experiments on CIFAR-10 and CIFAR-100, EDNAG achieves comparable performance on test accuracies compared to previous baselines.
-On Aircraft and Pets, EDNAG surpasses previous baselines with improvements of 2.23% and 10.45% on test accuracies, respectively.
+Each cell in NAS-Bench-201 includes 6 layers and 5 predefined operations for each layer.
+Within the NAS-BENCH-201 search spacee, we perform conditional generation of architectures across four task datasets and subsequently train the top-performing architecture to validate their performance, as shown in Table 1.
+In experiments on CIFAR-10 and CIFAR-100, EDNAG achieves comparable performance on test accuracies compared to previous methods.
+In Aircraft and Pets, EDNAG surpasses previous baselines with improvements of 2.23% and 10.45% on test accuracies, respectively.
 Across diverse task datasets, EDNAG outperforms all previous NAS and NAG methods, achieving state-of-the-art (SOTA) performance and showcasing exceptional capability in generating optimal architectures.
@@
 For CIFAR-10 and CIFAR-100, the final accuracies are obtained from the NAS-Bench-201 benchmarks Dong and Yang (2020) by querying only the top-1 generated architecture.
-For Aircraft and Pets, only the top-1 architecture requires full training in EDNAG, enhancing validation efficiency by 3$\times$ and 2$\times$ compared to DiffusionNAG.
-Results illustrate that the top-1 neural architecture identified by our neural predictor attains optimal accuracy after training and testing, demonstrating its remarkable performance, achieved by the application of coreset selection prior to dataset encoding.
+For Aircraft and Pets, only the top-1 architecture requires full training in EDNAG, enhancing the validation efficiency of 3$\times$ and 2$\times$ compared to DiffusionNAG.
+The results illustrate that the top-1 neural architecture identified by our neural predictor attains optimal accuracy after training and testing, demonstrating its remarkable performance, achieved by the application of coreset selection before dataset encoding.
 Unlike the random sampling method used in DiffusionNAG, this approach significantly enhances the stability and reliability of the predicted accuracies.
+4.1.3 Experiments on NAS-Bench-101 and NAS-Bench-301 Search Space
+Table 2: Experiments on NAS-BENCH-101 and NAS-Bench-301 search space. Accuracies (%) are reported over 3 runs on CIFAR-10 with 95% confidence intervals.
+Method NAS-Bench-101 NAS-Bench-301
+RS Bergstra and Bengio (2012) 93.59$\pm$0.42 94.40$\pm$0.08
+AG-Net Xiao and Wang (2024) 94.18$\pm$0.07 94.45$\pm$0.12
+SRLNAS Cassimon et al. (2025) 93.86$\pm$0.14 94.53$\pm$0.04
+DiNAS Asthana et al. (2024) 94.27$\pm$0.20 94.54$\pm$0.15
+EDNAG (Ours) 94.46$\pm$0.01 94.64$\pm$0.05
+Global Best 94.47 94.69
+NAS-Bench-101 is a cell-based search space encompassing $4.23\times 10^{5}$ architectures while NAS-Bench-301 is a DARTS Liu et al. (2018)-like search space consisting over $10^{10}$ architectures.
+We conduct experiments on NAS-Bench-101 and NAS-Bench-301 for image classification task on CIFAR-10, displayed in Table 2.
+EDNAG outperforms previous methods, particularly surpassing previous SOTA methods, by improving accuracy by 0.2% and 0.1% respectively, which highlights the strong capability of EDNAG in generating task-optimal architectures.
 4.1.4 Experiments on TransNASBench-101 Search Space
-Table 3: Experiments on TransNASBench-101 Search Space across seven tasks.
+Table 3: Experiments on TransNASBench-101 across seven tasks.
 Metrics are reported with 95% confidence intervals over 3 runs.
@@
 CATCH Chen et al. (2020) 45.27$\pm$0.5 54.38$\pm$0.2 56.13$\pm$0.7 56.99$\pm$0.6 25.38$\pm$0.4 60.70$\pm$0.7 -
-EDNAG (Ours) 46.32$\pm$0.0 54.94$\pm$0.0 57.72$\pm$0.0 59.62$\pm$0.0 26.27$\pm$0.0 58.71$\pm$0.0 95.37$\pm$0.0
+GEA Lopes et al. (2024) 45.98$\pm$0.2 54.85$\pm$0.1 57.11$\pm$0.3 58.33$\pm$1.0 25.95$\pm$0.2 59.93$\pm$0.5 94.96$\pm$0.2
+EDNAG (Ours) 46.28$\pm$0.04 54.91$\pm$0.03 57.63$\pm$0.09 59.26$\pm$0.36 26.19$\pm$0.08 59.01$\pm$0.30 95.19$\pm$0.18
 Global Best 46.32 54.94 57.72 59.62 26.27 58.71 95.37
@@
 CATCH Chen et al. (2020) 47.29$\pm$0.3 56.49$\pm$0.3 70.36$\pm$3.0 60.85$\pm$0.7 28.71$\pm$0.4 59.37$\pm$0.6 -
-EDNAG (Ours) 47.96$\pm$0.0 57.48$\pm$0.0 76.88$\pm$0.0 64.35$\pm$0.0 29.66$\pm$0.0 55.68$\pm$0.0 97.02$\pm$0.0
+EDNAG (Ours) 47.69$\pm$0.27 57.29$\pm$0.19 75.90$\pm$0.98 63.50$\pm$0.85 29.60$\pm$0.06 56.20$\pm$0.52 96.98$\pm$0.04
 Global Best 47.96 57.48 76.88 64.35 29.66 55.68 97.02
 TransNASBench-101 extends downstream tasks beyond basic image classification, encompassing a diverse set of 7 tasks: Object Classification, Scene Classification, Autoencoding, Surface Normal Estimation, Semantic Segmentation, Room Layout Prediction, and Jigsaw Puzzle Solving.
-Moreover, it defines two distinct types of search spaces: a cell-based micro search space and a stack-based macro skeleton search space, which encompass 4,096 and 3,256 unique architectures, respectively.
+Moreover, it defines two types of search spaces: a cell-based micro search space and a stack-based macro skeleton search space, which encompass 4,096 and 3,256 unique architectures, respectively.
 We conduct experiments in both micro-level and macro-level search spaces, as presented in Table 3.
-Across both search spaces, EDNAG exhibits significant performance improvements compared to previous baselines and consistently achieves globally optimal results across all downstream tasks while maintaining remarkable stability across three independent runs.
+Across both search spaces, EDNAG demonstrates substantial performance enhancements of up to 7.02% over prior baselines, consistently approaching global optimality across all downstream tasks while maintaining stability across three independent runs.
 These findings not only demonstrate the effectiveness of the FD strategy in improving global optimization capabilities for generative models, but also highlight the transferability of EDNAG for cross-task neural architecture generation within various downstream scenarios.
@@
 EDNAG (Ours) 97.40$\pm$0.01 85.81$\pm$0.02 81.87$\pm$0.10 95.06$\pm$0.14
-Due to the relatively limited scale of the above search spaces, our study further extends to the larger and more complex MobileNetV3 search space to demonstrate the scalability of EDNAG.
+Due to the relatively limited scale of the above search spaces, our study further extends to a larger and more complex MobileNetV3 search space to demonstrate the scalability of EDNAG.
 MobileNetV3 is a large search space which encompasses over $2.17\times 10^{19}$ distinct neural architectures, making it more than $10^{15}$ times larger than the NAS-Bench-201 search space.
-MobileNetV3 offers five stages, each consisting of 2 to 4 building blocks, where each building block is defined by the kernel size of convolutional layers chosen from {3, 5, 7} and the expansion rate of output channels chosen from {3, 4, 6}.
-We conduct experiments on four distinct task datasets within the MobileNetV3 search space, as shown in Table 4.
-In CIFAR-10, CIFAR-100, Aircraft, and Pets, EDNAG demonstrates an improvement in accuracy of up to 0.49%.
+We conduct experiments on four distinct task datasets within the MobileNetV3 search space, as shown in Table 4, where EDNAG demonstrates an improvement in accuracy of up to 0.49%.
 This experiment not only demonstrates the superior ability of architecture generation for EDNAG but also confirms its adaptability and effectiveness in large-scale search spaces.
-This is attributed to the fact that when transitioning to different search spaces, DiffusionNAG necessitates retraining the score network within the diffusion model to accommodate the noise matrix of the new dimensionality, which is troublesome and time-consuming.
+This adaptability is attributed to the fact that when transitioning to different search spaces, DiffusionNAG necessitates retraining the score network within the diffusion model to accommodate the new dimensionality of noise matrix, which is troublesome and time-consuming.
 In contrast, as a training-free NAG method, EDNAG avoids the need for training or retraining, thereby enabling adaptation to new search spaces with negligible costs.
@@
 To deeply analyze their critical roles, we conducted a comparative analysis of the architecture generation process under three distinct conditions: naive diffusion models, diffusion models with the FD strategy, and diffusion models with the FD strategy and optimization strategy, as illustrated in Figure 4.
-Naive diffusion models often results in convergence to local optima, resulting in suboptimal architecture generation.
+Naive diffusion models often result in convergence to local optima, resulting in suboptimal architecture generation.
 The integration of the FD strategy significantly enhances the accurancy of generated architectures, facilitating the generation of globally optimal solutions.
@@
 Pets 38 263
-As the first network-free neural architecture generation method, EDNAG boasts remarkable characteristics, including minimal computational resource consumption and rapid conditional generation, etc.
+As the pioneering network-free NAG approach, EDNAG achieves minimal computational resource consumption and rapid conditional generation.
 To demonstrate its efficiency, we compare EDNAG with DiffusionNAG in terms of computational costs during both training and inference processes in the NAS-Bench-201 search space, as shown in Table 5.
@@
 Therefore, EDNAG not only enables neural architecture generation without time-consuming training processes but also significantly reduces computational cost during the inference process.
-This is attributed to the FD strategy during the generative process, which directs the denoising process without relying on a score network, thereby achieving remarkable execution efficiency.
+This is attributed to the FD strategy during the generation process, which directs the denoising process without relying on a score network, thereby achieving remarkable execution efficiency.
 In contrast, DiffusionNAG depends on a complex score network at each step of the generative process to predict the noise to be removed, which demands extensive and frequent network inference, significantly slowing down the generation process.
@@
 In the NAS-BENCH-201 search space and CIFAR-10 dataset, we generate 1000 neural architectures using three different methods and visualize their accuracy distributions.
-We generate 1000 neural architectures by EDNAG and DiffusionNAG, then we display their distribution in Figure 6.
+We generate 1000 neural architectures by EDNAG and DiffusionNAG, then display their distribution in Figure 5.
 Compared to DiffusionNAG, EDNAG demonstrates a higher proportion of high-performing architectures.
-This demonstrates that the architectures generated by EDNAG exhibit a denser distribution within high-accuracy subspaces, reflecting a marked improvement in the quality of architecture generation.
-This advancement is achieved through the powerful conditional guidance and superior optimization ability by combining evolutionary algorithms, which contrasts sharply with the straightforward use of naive diffusion models in DiffusionNAG.
+This shows that the architectures generated by EDNAG exhibit a denser distribution within high-accuracy subspaces, reflecting a marked improvement in the quality of architecture generation.
+This advancement is achieved through the powerful conditional guidance and superior optimization ability by combining evolutionary algorithms, which guides architecture samples into a more stable and optimal subspace.
 4.3.3 Analysis of Neural Predictor
-Figure 7: Performance analysis of neural predictor.
-We predict the accuracy of architectures for CIFAR-10 and CIFAR-100, and display the Pearson correlation coefficient between predicted and actual accuracies.
-Figure 8: Stability analysis of neural predictor.
-With neural predictors in DiffusionNAG and EDNAG, we assessed the accuracy of identical neural architectures on consistent task datasets across 100 repeated runs.
-We conduct experiments within the NAS-BENCH-201 search space and report Pearson correlation coefficients between predicted and actual accuracies, as illustrated in Figure 7.
-On CIFAR-10, the predictor of DiffusionNAG suffers from biased predictions due to random sampling, resulting in a lower correlation coefficient, while EDNAG achieves a 12.09% improvement.
-On CIFAR-100, the predictor of EDNAG achieves a 7.29% enhancement compared to DiffusionNAG.
-This experiment reveals that the neural predictor of EDNAG, which utilizes K-Means clustering to select coreset images for dataset encoding, aligns more closely with real architectural accuracy.
-This approach effectively circumvents the limitations of random sampling in DiffusionNAG, which fails to preserve the uniformity of the target dataset distribution, resulting in biased predictions.
-Moreover, during the denoising iterations, identical architectures in successive generations ($x_{t}$ and $x_{t-1}$) necessitate multiple evaluations of the same architecture.
-Our experiments, shown in Figure 8, evaluate the same neural architecture over 100 repeated runs.
-The results demonstrate significant fluctuations in the predicted accuracy of DiffusionNAG due to random image sampling.
-In contrast, EDNAG demonstrates consistent predicted accuracy, effectively avoiding random perturbations in fitness values and further enhancing the stability of the whole generation process.
+Figure 6: Pearson correlation coefficients analysis of neural predictor.
+Figure 7: Stability analysis of neural predictor over 100 repeated runs.
+We report Pearson correlation coefficients between predicted and actual accuracies in NAS-BENCH-201, as illustrated in Figure 7.
+On CIFAR-10 and CIFAR-100, the predictor of DiffusionNAG suffers from biased predictions due to random sampling, resulting in a lower correlation coefficient, while EDNAG achieves improvements of 12.09% and 7.29%, respectively, by K-Means clustering to select coreset images before dataset encoding.
+Moreover, during denoising iterations, identical architectures in successive generations ($x_{t}$ and $x_{t-1}$) require multiple evaluations of the same architecture.
+Our experiments in Figure 7, evaluate a same neural architecture over 100 repeated runs.
+The results show significant fluctuations in the predicted accuracy of DiffusionNAG due to random image sampling.
+In contrast, EDNAG demonstrates nearly consistent predicted accuracy with its variance reduced by 48$\times$, effectively avoiding random perturbations in fitness values and further enhancing the stability of the whole generation process.
 5 Conclusion
@@
 Extensive experiments demonstrate that EDNAG achieves SOTA performance in generating task-optimal architectures and surpasses existing baselines in generation efficiency and computational cost.
-We believe that EDNAG will significantly contribute to the development of NAS by reducing computational costs and accelerate its practical application.
+We believe that EDNAG will contribute significantly to the development of NAS by reducing computational costs and accelerating practical application.